OpenStudy (anonymous):

Someone Please Help? Write an equation in slope-intercept form of the line with an x-intercept of 6 and a y-intercept of 5. A. y = 6x + 5 B. y = -5/6 x + 5 C. y = -6x + 5 D. y = 5/6 x + 5

OpenStudy (anonymous):

Intercept form of line = x/a + y/b = 1 here a = 6 and b =5 so x/6 + y/5 = 1 or 5x + 6y = 30 or y = -5/6 x + 5